Now suppose that in response to a stimulus package, the debt-to income ratio rises to 104% which causes an increase in the interest rate to 7.7% and leaves the primary deficit-to-GDP ratio constant. The debt-to-GDP ratio will now increase by ___________.   


1
Expert's answer
2021-05-18T12:46:37-0400

Given that,

rate of growth (g) = 0.25% ,

  interest rate(i) = 7.7%,

  debt-to-income ratio (B/Y) = 1.04 (104%

  and a primary deficit = 1.3%.


Using the formula,

Change in debt-to-GDP ratio = Primary deficit + (Interest rate - Growth rate) x (Debt-to-income ratio)


"= 1.3 + (7.70 - 0.25 ) \times ( 1.04)
\\


= 1.3 + (7.45)(1.04)\\ = 1.3 + 7.748\\ = 9.048\\"

"= 9.05%" %